Suppose you observe a red line in the hydrogen emission spectrum. For this specific transition, you know the final energy state is n=2, what must the initial energy state be? O A) n=4 OB) n=1 O C) n=2 O D) n=3
Answer #1

Since it is emission, initial should be greater than final state

So, initial state should be greater than 2

lets assume ni = 3

Here photon will be emitted

1/lambda = R* (1/nf^2 - 1/ni^2)

R is Rydberg constant. R = 1.097*10^7

1/lambda = R* (1/nf^2 - 1/ni^2)

1/lambda = 1.097*10^7* (1/2^2 - 1/3^2)

lambda = 6.563*10^-7 m

lambda = 656 nm

lets assume ni=4

Here photon will be emitted

1/lambda = R* (1/nf^2 - 1/ni^2)

R is Rydberg constant. R = 1.097*10^7

1/lambda = R* (1/nf^2 - 1/ni^2)

1/lambda = 1.097*10^7* (1/2^2 - 1/4^2)

lambda = 4.862*10^-7 m

lambda = 486 nm

lambda for red is around 656 nm

So, 1st assumption is correct

Answer: D